OpenStudy (anonymous):

Match the equation with its graph. –4x – 2y = 8

OpenStudy (johnweldon1993):

Well we can make this easier..by turning this equation into slope intercept form The form is y = mx + b where 'm' is slope and 'b' is y-intercept...so we need to solve your equation for 'y' -4x - 2y = 8 Start by adding 4x to both sides -2y = 4x + 8 and now divide EVERYTHING by -2 to isolate 'y' y = -2x - 4 So which of YOUR graphs...has slope -2 and y-intercept -4?

OpenStudy (whpalmer4):

find the x and y intercepts: -4(0)-2y=8 y = -4 (0,-4) -4x-2(0) = 8 -4x=8 x=-2 (-2,0) find the graph that goes through those two points.

OpenStudy (whpalmer4):

come on, you need to be able to read a graph! look for the one where the line goes through the points (-2,0) and (0,-4). Or look for the one that falls 2 points for every point you go to the right, and intersects the y-axis at -4.

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

What are the slope and y-intercept of the graph of the given equation? y = –4x + 2

OpenStudy (whpalmer4):

put it in slope-intercept form: \[y=mx+b\]then \(m\) is slope and \(b\) is y-intercept